For an economy the following functions have been given:
C = 100 + 0.8Y
I = 120 – 5r
Ms = 120
Md = 0.2Y – 5r
Calculate the following:
1 IS equation
2. LM equation
3. Equilibrium level of income
4. Equilibrium level of interest rate.
1
Expert's answer
2019-08-14T09:26:32-0400

IS Equation

Y = AD = C + I

Y = 100 + 0.8 Y + 120 - 5r


LM Equation

Md=Ms


120=0.2y -5r

5r = 0.2y -120

r = 0.04y + 24


Equilibrium Level of income

Y = 100 + 0.8y +120 - 5(0.04y + 24)

Y = 250


Equilibrium Level of Interest rate

r = 0.04(250) + 24

r = 34%